# A variational discretization method for mean curvature flows by the Onsager principle

### v1

@@ 8 Acknowledgement
The right-hand term $\tilde{G}=(\tilde{\vec{g}}_{1}^{T},\tilde{\vec{g}}_{2}^{T},...,\tilde{\vec{g}}_{n}^{T},\tilde{g}_{n+1})^{T}$, where $\tilde{\vec{g}}_{i}(1\leq i\leq n)$ denote two dimensional column vectors and $\tilde{g}_{n+1}$ denotes a scalar, respectively. They are calculated as
$\tilde{\vec{g}}_{1}=(-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial x_{1}^{(1)}},0)^{T}=(\gamma[\frac{x_{2}^{(1)}(t)-x_{1}^{(1)}(t)}{|\vec{x}_{2}(t)-\vec{x}_{1}(t)|}+\cos\theta_{Y}],0)^{T}\quad for\quad i=1$ (108)
$\tilde{\vec{g}}_{i}=-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial\vec{x}_{i}}=\gamma[\frac{\vec{x}_{i+1}(t)-\vec{x}_{i}(t)}{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}-\frac{\vec{x}_{i}(t)-\vec{x}_{i-1}(t)}{|\vec{x}_{i}(t)-\vec{x}_{i-1}(t)|}]\quad for\quad 2\leq i\leq n-1.$
$\tilde{\vec{g}}_{n}=(-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial x_{n}^{(1)}},0)^{T}=(-\gamma[\frac{x_{n}^{(1)}(t)-x_{n-1}^{(1)}(t)}{|\vec{x}_{n}(t)-\vec{x}_{n-1}(t)|}+\cos\theta_{Y}],0)^{T}\quad for\quad i=n$
$\tilde{g}_{n+1}=0\quad for\quad i=n+1.$
[…]

### v3

@@ Acknowledgments
The right-hand term $\tilde{G}=(\tilde{\vec{g}}_{1}^{\top},\tilde{\vec{g}}_{2}^{\top},...,\tilde{\vec{g}}_{n}^{\top},\tilde{g}_{n+1})^{\top}$, where $\tilde{\vec{g}}_{i}\ (1\leq i\leq n)$ denote two dimensional column vectors and $\tilde{g}_{n+1}$ denotes a scalar, respectively. They are calculated as
$\tilde{\vec{g}}_{i}=-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial\vec{x}_{i}}=\gamma[\frac{\vec{x}_{i+1}(t)-\vec{x}_{i}(t)}{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}-\frac{\vec{x}_{i}(t)-\vec{x}_{i-1}(t)}{|\vec{x}_{i}(t)-\vec{x}_{i-1}(t)|}],$ $2\leq i\leq n-1,$
$\tilde{\vec{g}}_{1}=(-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial x_{1}^{(1)}},0)^{\top}=(\gamma[\frac{x_{2}^{(1)}(t)-x_{1}^{(1)}(t)}{|\vec{x}_{2}(t)-\vec{x}_{1}(t)|}+\cos\theta_{Y}],0)^{\top},$
$\tilde{\vec{g}}_{n}=(-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial x_{n}^{(1)}},0)^{\top}=(-\gamma[\frac{x_{n}^{(1)}(t)-x_{n-1}^{(1)}(t)}{|\vec{x}_{n}(t)-\vec{x}_{n-1}(t)|}+\cos\theta_{Y}],0)^{\top},$
$\tilde{g}_{n+1}=0,$ $i=n+1.$